OpenStudy (anonymous):

8x + 13y = 2 y = x- 2 Solve the system of equations A.(4/3, -2/3) B.(4/21, -38/21) C.no solution

OpenStudy (anonymous):

okay you can solve this system of equations by two methods: elimination and substitution. Because y is given (in the second equation), substitution method would be the easiest.

OpenStudy (anonymous):

so plug in y=x-2 in place of y in the first equation.

OpenStudy (anonymous):

can you answer ?

OpenStudy (anonymous):

okay... so, 8x+13(x-2) = 2 8x+13x-26=2 21x-26=2 21x=2+26 21x=28 x=28/21 x=4/3

OpenStudy (anonymous):

alright :) thanks for helping i will give you medal ;)

OpenStudy (anonymous):

now that you have x, plug in x in either of the equations to find out y. since eq. 2 is easier, i would do 4/3 - 2 = -2/3 is your y.